# CLEF 2025 – Conference and Labs of the Evaluation Forum, September 9–12, 2025, Madrid, Spain

### v1

@@ Abstract
The large language models (LLMs) are able to generate high-quality texts in multiple languages. Such texts are often not recognizable by humans as generated, and therefore present a potential of LLMs for misuse (e.g., plagiarism, spams, disinformation spreading). An automated detection is able to assist humans to indicate the machine-generated texts; however, its robustness to out-of-distribution data is still challenging. This notebook describes our mdok approach in robust detection, based on fine-tuning smaller LLMs for text classification. It is applied to both subtasks of Voight-Kampff Generative AI Detection 2025, providing remarkable performance in binary detection as well as in multiclass (1st rank) classification of various cases of human-AI collaboration.
@@ 1 Introduction
The Voight-Kampff Generative AI Detection 2025 shared task, as a part of PAN lab [5] at the CLEF 2025 conference, addresses two challenges in the machine-generated text detection problem area in two subtasks. Subtask 1 is focused on classical binary detection task distinguishing between machine-authored (class 1) and human-authored (class 0) texts, which addresses also the robustness of detectors to obfuscation and other surprise data. Subtask 2 is focused on human-AI collaborative text classification, covering six classes of collaboration, namely class 0 representing fully human-written texts; class 1 representing human-written, then machine-polished texts, class 2 representing machine-written, then machine-humanized texts, class 3 representing human-initiated, then machine-continued texts, class 4 representing deeply-mixed text (where some parts are written by a human and some are generated by a machine), and class 5 representing machine-written, then human-edited texts.
In this notebook, we describe our mdok approach addressing both subtasks (two separate systems). We build on our previously developed robust LLM fine-tuning for sequence classification [6], but adjust the training process to the shared task training data and binary/multiclass problem formulation. We further explore the usage of the most recent LLMs, size of which ranging from 1B to 14B parameters. The submitted systems are based on Qwen3-4B and Qwen3-14B models [7].
@@ 2 Background
A multiclass machine-generated text classification is mostly researched in related authorship attribution task, identifying of the author (generator) of the text. The problem of different classes of AI-human collaboration can be approached by existing authorship attribution methods [14, 15], only slightly different from binary (two-class) methods.
@@ 3 System Overview
For the fine-tuning process, we have used the adjusted published script for QLoRA based robust fine-tuning of [6]. Besides of modification of training data mixture, we have used learning rate of $42e-$ 4, avoided gradient accumulation, and limited the training time to 3 epochs. Due to already incorporated weighted cross entropy for loss calculations, it was straightforward to extend it to multiclass classification by calculating weights based on the training data class distribution.
@@ 4.1 Subtask 1 Official Results
The official results of Subtask 1 have not been released before submitting the system description notebook.
@@ 4.2 Subtask 2 Official Results
Table 4: Official Leaderboard of Subtask 2 (for Teams Ourperforming the Baseline)
@@ 5 Conclusion
The robustness of machine-generated text detection methods to out-of-distribution data is still challenging. However, we can cope with this problem by better mixture of training data, as we have show in the proposed mdok approach for both binary and multiclass detection. The resulting detectors are very competitive (ranking 1st in multiclass detection of human-AI collaboration), outperforming all the baselines, and promising better generalization to out-of-distribution data. Further work is still needed to refine the process to achieve ideal performance.
@@ Acknowledgements.
This work was partially supported by the European Union NextGenerationEU through the Recovery and Resilience Plan for Slovakia under the project No. 09I01-03-V04-00059 and partially by LorAI – Low Resource Artificial Intelligence, a project funded by Horizon Europe under GA No.101136646. We acknowledge EuroHPC Joint Undertaking for awarding us access to Leonardo at CINECA, Italy.

### v2

# CLEF 2025 Working Notes, 9 – 12 September 2025, Madrid, Spain

@@ Abstract
The large language models (LLMs) are able to generate high-quality texts in multiple languages. Such texts are often not recognizable by humans as generated, and therefore present a potential of LLMs for misuse (e.g., plagiarism, spams, disinformation spreading). An automated detection is able to assist humans to indicate the machine-generated texts; however, its robustness to out-of-distribution data is still challenging. This notebook describes our mdok approach in robust detection, based on fine-tuning smaller LLMs for text classification. It is applied to both subtasks of Voight-Kampff Generative AI Detection 2025, providing remarkable performance (1st rank) in both, the binary detection as well as the multiclass classification of various cases of human-AI collaboration.
@@ 1 Introduction
The Voight-Kampff Generative AI Detection 2025 shared task [5], as a part of PAN lab [6] at the CLEF 2025 conference, addresses two challenges in the machine-generated text detection problem area in two subtasks. Subtask 1 is focused on classical binary detection task distinguishing between machine-authored (class 1) and human-authored (class 0) texts, which addresses also the robustness of detectors to obfuscation and other surprise data. Subtask 2 is focused on human-AI collaborative text classification, covering six classes of collaboration, namely class 0 representing fully human-written texts; class 1 representing human-written, then machine-polished texts, class 2 representing machine-written, then machine-humanized texts, class 3 representing human-initiated, then machine-continued texts, class 4 representing deeply-mixed text (where some parts are written by a human and some are generated by a machine), and class 5 representing machine-written, then human-edited texts.
In this notebook, we describe our mdok approach addressing both subtasks (two separate systems). We build on our previously developed robust LLM fine-tuning for sequence classification [7], but adjust the training process to the shared task training data and binary/multiclass problem formulation. We further explore the usage of the most recent LLMs, size of which ranging from 1B to 14B parameters. The submitted systems are based on Qwen3-4B and Qwen3-14B models [8]. For the replication purpose, we publish the source code of the mdok approach [^1] https://github.com/kinit-sk/mdok.
@@ 2 Background
A multiclass machine-generated text classification is mostly researched in related authorship attribution task, identifying the author (generator) of the text. The problem of different classes of AI-human collaboration can be approached by existing authorship attribution methods [15, 16], only slightly different from binary (two-class) methods.
@@ 3 System Overview
For the fine-tuning process, we have used the adjusted published script for QLoRA based robust fine-tuning of [7]. Besides of modification of training data mixture, we have used learning rate of 2e42e-4, avoided gradient accumulation, and limited the training time to 3 epochs. Due to already incorporated weighted cross entropy for loss calculations, it was straightforward to extend it to multiclass classification by calculating weights based on the training data class distribution.
@@
3.1 Homoglyph Attack
As mentioned, we have used a generic homoglyph attack of [17] to modify (in the pre-processing step) a small portion of training data (pseudo-randomly selected 10% of machine-generated texts) for Subtask 1 to increase its robustness to obfuscation. It uses the whole confusables table [^2] https://www.unicode.org/Public/security/8.0.0/confusables.txt to psudorandomly replace letters for their homoglyphs. The probability of a character to be replaced was set to 0.05 (i.e., about 5% of characters). It further integrates a pseudorandom insertion of visually unseenable zero-width-joiner character in the text (also with a probability of 0.05). We have used the random seed of 42. It has been previously shown that seeing such obfuscated data in the training (fine-tuning) effectively eliminates their highly negative effect on the detection.
3.2 MIX2k Dataset
For validation in out-of-distribution settings in Subtask 1, we have used MIX2k dataset, introduced by [7]. It contains 1,000 samples for each class (human and machine), pseudo-randomly sampled from 18 existing labeled datasets. Such data composition consists of data from 75 generators in 7 languages, thus being mostly out-of-distribution (in comparison to training data). Validation on such a dataset during fine-tuning procedure helps to select the model checkpoint, which has the best generalization to out-of-distribution data, effectively avoiding overfit to training data.
@@ 4.1 Subtask 1 Official Results
The official results are provided in Table 4, where Mean of the official metrics has been used for ranking. The results show that our approach outperforms all the others in all metrics except of ROC-AUC, where the system of one team outperformed our detector. The performances of the systems are close enough to yield a more efficient detection system, since we have relied on a rather large (14B of parameters) model, consuming more resources for inference. However, the mdok single-model detection system can still be more efficient than some more complex ensemble systems, relying on multiple inference rounds of multiple (although smaller) models. If the labeled test set will be released, an ablation study could pinpoint the strong parts of our system that make it the best performing and/or identify a smaller model to be used instead to provide a more efficient solution.
Table 4: Official Leaderboard of Subtask 1 (for Teams Outperforming the Best-performing Baseline)
Rank Team Name ROC-AUC Brier C@1 F1 F0.5u Mean
1 mdok (binary) 0.853 0.896 0.894 0.898 0.903 0.899
2 steely 0.842 0.879 0.877 0.865 0.881 0.880
3 nexus-interrogators 0.865 0.874 0.870 0.860 0.881 0.879
4 yangjlg 0.845 0.878 0.871 0.856 0.881 0.877
5 cnlp-nits-pp 0.825 0.873 0.873 0.854 0.882 0.874
6 unibuc-nlp 0.828 0.885 0.864 0.845 0.876 0.872
7 moadmoad 0.822 0.866 0.865 0.855 0.882 0.871
8 iimasnlp 0.838 0.868 0.856 0.851 0.877 0.869
9 bohan-li 0.848 0.858 0.852 0.847 0.870 0.866
10 advacheck 0.802 0.855 0.855 0.854 0.879 0.863
11 hello-world 0.838 0.871 0.836 0.827 0.862 0.856
@@ 4.2 Subtask 2 Official Results
Table 5: Official Leaderboard of Subtask 2 (for Teams Outperforming the Baseline)
@@ 5 Conclusion
The robustness of machine-generated text detection methods to out-of-distribution data is still challenging. However, we can cope with this problem by better mixture of training data, as we have show in the proposed mdok approach for both binary and multiclass detection. The resulting detectors are very competitive (ranking 1st in both subtasks), outperforming all the baselines, and promising better generalization to out-of-distribution data. Due to gold labels of the test data are not available yet, any ablation study pinpointing crucial parts of the systems and their specific effects was not possible. Further work is still needed to refine the process to achieve ideal performance.
@@ Acknowledgements.
This work was partially supported by the European Union NextGenerationEU through the Recovery and Resilience Plan for Slovakia under the project No. 09I01-03-V04-00059 and partially by LorAI – Low Resource Artificial Intelligence, a project funded by Horizon Europe under GA No.101136646. Computational resources. We acknowledge EuroHPC Joint Undertaking for awarding us access to Leonardo at CINECA, Italy.
Declaration on Generative AI
The author(s) have not employed any Generative AI tools.